OpenStudy (anonymous):

Which point lies on a circle that is centered at A(3, 3) and passes through B(6, 5)? C(1, 6) D(6, 0) E(0, 3) F(3, -1) G(3, 6)

OpenStudy (anonymous):

would you use mid point formula

OpenStudy (sheraz12345):

Yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Why do you need the mid-point formula?

OpenStudy (sheraz12345):

@eliassaab because we have the mid point given(A).. So we will find that option which will give same midpoint with B

OpenStudy (anonymous):

The equation of the circle is \[ (x-3)^2+(y-3)^2=13 \\ (1-3)^2 +(6-3)^2=4+9 =13 \] Hence (1,6) is on the circle.

OpenStudy (anonymous):

13 is calculated as the square distance from the given point to the center of the circle.

OpenStudy (sheraz12345):

Cant remember the formula for radius .. CAN U tell me @eliassaab

OpenStudy (anonymous):

Do you know how to find the distance between two points?